# GraphIC: A Graph-Based In-Context Example Retrieval Model for Multi-Step Reasoning

### v2

@@ Abstract
GraphIC: A Graph-Based In-Context Example Retrieval Model for Multi-Step Reasoning
Jiale Fu1, Yaqing Wang2,*, Simeng Han3, Jiaming Fan1, Xu Yang1,* 1Southeast University, 2Beijing Institute of Mathematical Sciences and Applications, 3Yale University, *Corresponding author.
@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020) allows large language models (LLMs) to adapt to new tasks by incorporating a few demonstration examples within the input prompt, without updating model parameters. However, studies reveal that ICL performance heavily depends on the quality of selected in-context examples (ICEs) Zhao et al. (2021); Liu et al. (2022), motivating extensive research on ICE selection strategies. Current approaches Liu et al. (2022); Rubin et al. (2022) typically use text embeddings to measure semantic similarity between queries and candidate examples, achieving success in semantic-centric tasks like text classification and translation Agrawal et al. (2023).
However, these text-based methods face significant limitations in multi-step mathematical and logical reasoning tasks. This is because text embedding encodes a substantial amount of shallow semantic information, which is irrelevant to the underlying reasoning processes. This extraneous information introduces bias in the selection of ICEs An et al. (2023). As shown in Figure 1 (left), when solving a speed calculation problem, text-based methods may retrieve examples about distance/time calculations due to shallow semantic similarity, inducing incorrect reasoning paths (e.g., calculating distance instead of speed). This observation motivates our key insight: Effective ICE selection for reasoning tasks requires representations that explicitly model cognitive processes rather than textual surface forms.
Drawing from cognitive science Friston (2008) and graph-based reasoning works Besta et al. (2024); Yao et al. (2023), we propose thought graphs—directed node-attributed graphs where nodes represent reasoning steps and edges denote step dependencies (i.e., a child step can only proceed after the parent step is completed). This representation filters irrelevant semantics while preserving essential reasoning patterns. An example of a thought graph is shown in Figure 2 (a).
Figure 1: ICL with different ICE retrieval mechanisms. The left panel shows examples retrieved via BERT embedding Devlin et al. (2019), while the right panel displays examples retrieved via GraphIC. Semantically related terms are highlighted in blue, and quantities or variables needing resolution are highlighted in green.
@@
Through comprehensive evaluations across mathematical reasoning, code generation, and logical reasoning tasks, GraphIC demonstrates superior performance over 10 baseline methods including both training-free and training-based approaches. To sum up, our key contributions are: (1) A representaion.We introduce a novel graph-based representation, called thought graph, to model multi-step reasoning processes. This representation effectively filters out irrelevant shallow semantic information while preserving the essential reasoning steps. (2) A similarity metric.We introduce a similarity metric for thought graphs tailored to multi-step reasoning tasks that capture the sequential nature of the steps and the asymmetry between examples. (3) Empirical validation.Our experimental results indicate that GraphIC, despite being a training-free model, outperforms both training-free and training-based models across various multi-step reasoning tasks.
@@ 2 Related Work
Training-free approaches are generally divided into two types: (i) those that use heuristic criteria such as similarity Liu et al. (2022); Hu et al. (2022), diversity Cho et al. (2023); Zhang et al. (2022b); Levy et al. (2023); Hongjin et al. (2022); Zhang et al. (2023), complexity Fu et al. (2022), or combinations of these Agrawal et al. (2023); Tonglet et al. (2023); Gupta et al. (2023) to select in-context examples (ICEs); (ii) those that leverage feedback from LLMs, such as probability distributions Wu et al. (2023); Nguyen and Wong (2023); Li and Qiu (2023); Yang et al. (2023), perplexity Gonen et al. (2023), or the model’s generated output An et al. (2023) to guide the selection process. Training-free approaches eliminate the computational and time costs of model training, but their simpler architecture often leads to lower performance than training-based methods.
Training-based methods are generally divided into two types. The first learns to select individual examples and then extends this to $k$-shot scenarios Rubin et al. (2022); Xiong et al. (2024); Gupta et al. (2024). The second models the selection of a group of examples as a whole Ye et al. (2023); Wang et al. (2023); Zhang et al. (2022a); Scarlatos and Lan (2023); Lu et al. (2022); Peng et al. (2023); Xu et al. (2024). Training-based approaches often deliver better performance, but they are computationally expensive and time-consuming.
@@ 3.1 Thought Graphs
[…]
@@
Due to the inherent differences between tasks, the text used to compute embedding varies slightly. Specifically, for mathematical reasoning, code generation, and logical reasoning tasks, we use mathematical operations, code snippets, and intermediate conclusions, respectively, as input texts. This is because, in mathematical reasoning tasks, a reasoning step can be represented by "which mathematical operations were performed on which variables"; in code generation tasks, it corresponds to "what code was executed"; and in logical reasoning tasks, it reflects "which existing conclusions were used to derive new conclusions". Examples of thought graphs for the four benchmark datasets are provided in Appendix A.1.
@@
Construction of Thought Graphs. For candidate examples, which include ground-truth natural language reasoning processes (i.e., chain-of-thought answers), we employ the following method to generate thought graphs. First, we use an LLM to formalize the reasoning process for each example. This formalized representation is called the Formalized Reasoning Representation (FRR). FRR filters irrelevant shallow semantics and extracts the key information from the reasoning process. Moreover, FRR is designed to be easily converted to a graph by applying rule-based parsing. The prompts used for generating FRRs and the pseudocode for parsing FRRs are shown in Appendix A.2.
@@
Figure 2 (b) illustrates the FRR of a thought graph. Each line in the FRR follows the format A = [B](C), where A represents the name of the vertex, denoting the vertex itself. B represents the label of the vertex, which will be used to compute its embedding. C denotes the parent vertex of the given vertex. For example, in the FRR shown in Figure 2 (b), by parsing the first line, we create node_1, set its label to "add", and add edges from "20" and "0.3" to node_1. The remaining lines are parsed similarly to the first one. Note that in Figure 2 (a), the vertices representing numbers are indicated by dashed lines, which signifies that these vertices do not represent specific operations and will be removed once the entire graph is constructed.
@@
For queries, since they lack ground-truth reasoning processes, we first use an LLM to generate pseudo natural language reasoning processes. We then follow the same procedure as with candidate examples to create thought graphs. Note that it is unnecessary to ensure the correctness of the generated reasoning process, as it is only used to aid in the retrieval of ICEs and is not part of the final answer. In fact, when a generated answer is incorrect, it is often not entirely wrong, but partially correct, which still reflects the reasoning process and can be helpful for retrieving examples. This ensures that GraphIC is robust to inaccuracies in the generated reasoning process. Even when the reasoning process is incorrect, GraphIC still maintains high accuracy. A detailed discussion and related experiments are provided in Appendix A.3.
@@ 3.2 Similarity Measure for Thought Graphs
For instance, consider a thought graph $G$ with two vertices: $v_{1}$ with attribute $x_{1}=\operatorname{Emb(\text{"add"})}$ and $v_{2}$ with $x_{2}=\operatorname{Emb(\text{"multiply"})}$, connected by an edge $v_{1}\to v_{2}$. In this setup, $W$ learns to map $x_{1}$ ("add") to $x_{2}$ ("multiply"), indicating that "multiply" typically follows "add" in $G$. When this learned mapping is applied to another graph $G^{\prime}$, a high value of $\mathcal{A}(W;G^{\prime})$ suggests that $G^{\prime}$ also contains frequent "add" $\to$ "multiply" sequences, reflecting similar problem-solving logic. Therefore, $\mathcal{A}(W;G^{\prime})$ quantifies how well the reasoning pattern of $G$ applies to $G^{\prime}$. A higher $\mathcal{A}(W;G^{\prime})$ indicates an alignment between the two graphs’ transitions, indicating that $G^{\prime}$ follows a similar problem-solving trajectory. This metric, therefore, provides an effective measure of the similarity in reasoning between $G$ and $G^{\prime}$, making it suitable for retrieving ICEs that share compatible problem-solving logic.
@@
where $Z^{(h)}$ represents node information states at iteration $h$; $Z^{(0)}$ is initialized with node attributes $X$; $I$ is identity matrix; $\tilde{A}=D_{A}^{-\frac{1}{2}}AD_{A}^{-\frac{1}{2}}$, $D_{A}=\operatorname{diag}(\operatorname{deg}_{\operatorname{out}}(v_{1}),\dots,\operatorname{deg}_{\operatorname{out}}(v_{1}))$; $\tilde{B}=D_{B}^{-\frac{1}{2}}BD_{B}^{-\frac{1}{2}}$, and $B$ is traceback matrix enabling backtracking to root nodes, defined as $B_{ij}=1$ if $\text{deg}_{\text{in}}(v_{j})=0$ and $\text{deg}_{\text{in}}(v_{i})>0$, otherwise $B_{ij}=0$, with $\text{deg}_{\text{in}}(v_{j})$ representing the in-degree of node $v_{j}$. A visual example of matrix $B$ is presented in Appendix A.4. $\lambda$ is a balancing hyperparameter ($0\leq\lambda\leq 1$).
@@
Both theoretical analysis and empirical evidence, presented in Appendix A.5, demonstrate that the rank-1 assumption has minimal impact on solving the optimization problem.
Furthermore, with the rank-1 assumption, we can obtain the closed-form solution of Equation (1), as shown in Equation (4), which greatly reduces the computational complexity of solving the optimization problem. The proof is shown in Appendix A.6
@@
$\text{where}\quad U,\Sigma,V=\operatorname{SVD}(X^{\top}Z).$
@@
Secondly, our proposed similarity is asymmetric. Specifically, for two thought graphs $G_{1}$ and $G_{2}$, $s(G_{1},G_{2})\neq s(G_{2},G_{1})$. This asymmetry reflects real-world scenarios. For instance, as illustrated in Appendix A.7, mathematical problem A might be a subproblem of problem B. In such a case, referencing B can help resolve A, but referencing A does not necessarily resolve B.
@@ 4.1 Experimental Setup
Datasets and Evaluations. We conduct experiments across four multi-step reasoning tasks: mathematical reasoning (GSM8K Cobbe et al. (2021) and AQUA Ling et al. (2017)), code generation (MBPP Austin et al. (2021)), and logical reasoning (ProofWriter Tafjord et al. (2021)). Model performance is measured by answer accuracy for GSM8K, AQUA, and ProofWriter, and by the pass@1 metric Chen et al. (2021) for MBPP.
Models and Hyper-parameters. We use both an open-source and a closed-source model. For the open-source model, we select Llama-3.1-8B-Instruct (hereafter referred to as Llama-3), one of the most advanced 7B-level models. For the closed-source model, we choose GPT-4o-mini, balancing performance and testing costs. Unless explicitly mentioned otherwise, all evaluations use an 8-shot setting, which is the most common setting in chain-of-thought scenarios. We set the temperature to 1e-5 to minimize randomness, ensuring consistency across generations, and use 3 iterations in Equation (2) with $\lambda$ values from $\{0,0.1,0.2,0.3\}$ (see Appendix 5 for specific details). More experiment details are in Appendix B.1.
@@ 4.2 Baselines
We compare GraphIC against six training-free retrieval methods spanning random, similarity-based, diversity-based, and complexity-based approaches, including: (1) Randomrandomly selects $k$ unique ICEs from the candidate set; (2) BM25Robertson et al. (2009)selects the top $k$ examples based on BM25 scoring; (3) BERTDevlin et al. (2019)is a dense retriever using cosine similarity with BERT-base-uncased embeddings; (4) Complex-CoTFu et al. (2022)selects $k$ examples based on complexity, quantified by newline characters; (5) Auto-CoTZhang et al. (2022b)clusters candidates and selects the closests to each cluster center; and (6) Skill-kNNAn et al. (2023)prompts LLM to generate task-relevant skills for query and candidates, followed by dense retrieval.
We also compare with four training-based methods, which encompass both single-example and combination-example retrieval strategies, including: (1) EPRRubin et al. (2022)is trained to retrieve the single most relevant ICE, with top $k$ examples being selected during inference; (2) CEILYe et al. (2023)uses determinantal point processes to select ICEs balancing similarity and diversity; (3) DQ-LoReXiong et al. (2024)uses dual queries and low-rank approximation re-ranking to identify ICEs; and (4) GistScoreGupta et al. (2024)encodes task-specific information into gist tokens for selecting ICEs.
@@ 4.3 Main Results
Llama-3.1 -8B-Instruct Random 78.86 (0.87) 53.15 (1.85) 57.72 (1.06) 76.10 (2.45) 66.46 (0.84)

### v4

@@ Abstract
Codehttps://github.com/jialefu/GraphIC
@@ 1 Introduction
In-context learning (ICL) (NEURIPS2020_1457c0d6; peng2024live; jiang2025mimic) allows large language models (LLMs) to adapt to new tasks by incorporating a few demonstration examples within the input prompt, without updating model parameters. However, studies reveal that ICL performance heavily depends on the quality of selected in-context examples (ICEs) (zhao2021calibrate; yang2023exploring; li2024configure), motivating extensive research on ICE selection strategies. Current approaches (liu2022makes; rubin2022learning) typically use text embeddings to measure semantic similarity between queries and candidate examples, achieving success in semantic-centric tasks like text classification and translation (agrawal2023context).
However, these text-based methods face significant limitations in multi-step mathematical and logical reasoning tasks. This is because text embedding encodes a substantial amount of shallow semantic information, which is irrelevant to the underlying reasoning processes. This extraneous information introduces bias in the selection of ICEs (an2023skill). As shown in Figure 1 (left), when solving a speed calculation problem, text-based methods may retrieve examples about distance/time calculations due to shallow semantic similarity, inducing incorrect reasoning paths (e.g., calculating distance instead of speed). This observation motivates our key insight: Effective ICE selection for reasoning tasks requires representations that explicitly model cognitive processes rather than textual surface forms.
Drawing from cognitive science (friston2008hierarchical) and graph-based reasoning works (besta2024graph; yao2024tree), we propose thought graphs—directed node-attributed graphs where nodes represent reasoning steps and edges denote step dependencies (i.e., a child step can only proceed after the parent step is completed). This representation filters irrelevant semantics while preserving essential reasoning patterns. An example of a thought graph is shown in Figure 2 (a).
Figure 1: ICL with different ICE retrieval mechanisms. The left panel shows examples retrieved via BERT embedding (devlin-etal-2019-bert), while the right panel displays examples retrieved via GraphIC. Semantically related terms are highlighted in blue, and quantities or variables needing resolution are highlighted in green.
@@
Through comprehensive evaluations across mathematical reasoning, code generation, and logical reasoning tasks, GraphIC demonstrates superior performance over 10 baseline methods including both training-free and training-based approaches. To sum up, our key contributions are:
  1. 1.
    A representaion. We introduce a novel graph-based representation, called thought graph, to model multi-step reasoning processes. This representation effectively filters out irrelevant shallow semantic information while preserving the essential reasoning steps.
  2. 2.
    A similarity metric. We introduce a similarity metric for thought graphs tailored to multi-step reasoning tasks that capture the sequential nature of the steps and the asymmetry between examples.
  3. 3.
    Empirical validation. Our experimental results indicate that GraphIC, despite being a training-free model, outperforms both training-free and training-based models across various multi-step reasoning tasks.
@@ 2 Related Work
Training-free approaches are generally divided into two types: (i) those that use heuristic criteria such as similarity (liu2022makes; hu2022context), diversity (cho2023prompt; zhangautomatic; levy2023diverse; hongjin2022selective; zhang2023ideal), complexity (fu2022complexity), or combinations of these (agrawal2023context; tonglet2023seer; gupta2023coverage) to select in-context examples (ICEs); (ii) those that leverage feedback from LLMs, such as probability distributions (wu2023self; nguyen2023context; li2023finding; yang2023representative), perplexity (gonen2023demystifying), or the model’s generated output (an2023skill) to guide the selection process. Training-free approaches eliminate the computational and time costs of model training, but their simpler architecture often leads to lower performance than training-based methods.
Training-based methods are generally divided into two types. The first learns to select individual examples and then extends this to $k$-shot scenarios (rubin2022learning; xiong2024dqlore; guptagistscore). The second models the selection of a group of examples as a whole (ye2023compositional; wang2024large; zhang2022active; scarlatos2023reticl; lu2022dynamic; xureprompting; yang2024lever). Training-based approaches often deliver better performance, but they are computationally expensive and time-consuming.
@@ 3.1 Thought Graphs
As illustrated in Figure 2 (a), the text corresponding to each vertex is ”divide”, ”add”, ”minus”, and ”multiply”. Therefore, the attributes of each vertex are $\mathrm{Emb}(\text{"divide"})$, $\mathrm{Emb}(\text{"add"})$, $\mathrm{Emb}(\text{"minus"})$, and $\mathrm{Emb}(\text{"multiply"})$, representing four base operations in a mathematical reasoning task. Here, $\mathrm{Emb}(\cdot)$ represents calculating the embedding of a given text using an embedding model, such as BERT. In this graph, an edge from the vertex labeled ”add” to the vertex labeled ”multiply” indicates that addition is performed first, and its result is then used in multiplication.
@@
Due to the inherent differences between tasks, the text used to compute embedding varies slightly. Specifically, for mathematical reasoning, code generation, and logical reasoning tasks, we use mathematical operations, code snippets, and intermediate conclusions, respectively, as input texts. This is because, in mathematical reasoning tasks, a reasoning step can be represented by ”which mathematical operations were performed on which variables”; in code generation tasks, it corresponds to ”what code was executed”; and in logical reasoning tasks, it reflects ”which existing conclusions were used to derive new conclusions”. Examples of thought graphs for the four benchmark datasets are provided in the supplementary material.
@@
Construction of Thought Graphs. For candidate examples, which include ground-truth natural language reasoning processes (i.e., chain-of-thought answers), we employ the following method to generate thought graphs. First, we use an LLM to formalize the reasoning process for each example. This formalized representation is called the Formalized Reasoning Representation (FRR). FRR filters irrelevant shallow semantics and extracts the key information from the reasoning process. Moreover, FRR is designed to be easily converted to a graph by applying rule-based parsing. The prompts used for generating FRRs and the pseudocode for parsing FRRs are shown in the supplementary material.
@@
Figure 2 (b) illustrates the FRR of a thought graph. Each line in the FRR follows the format A = [B](C), where A represents the name of the vertex, denoting the vertex itself. B represents the label of the vertex, which will be used to compute its embedding. C denotes the parent vertex of the given vertex. For example, in the FRR shown in Figure 2 (b), by parsing the first line, we create node_1, set its label to ”add”, and add edges from ”20” and ”0.3” to node_1. The remaining lines are parsed similarly to the first one. Note that in Figure 2 (a), the vertices representing numbers are indicated by dashed lines, which signifies that these vertices do not represent specific operations and will be removed once the entire graph is constructed.
@@
For queries, since they lack ground-truth reasoning processes, we first use an LLM to generate pseudo natural language reasoning processes. We then follow the same procedure as with candidate examples to create thought graphs. Note that it is unnecessary to ensure the correctness of the generated reasoning process, as it is only used to aid in the retrieval of ICEs and is not part of the final answer. In fact, when a generated answer is incorrect, it is often not entirely wrong, but partially correct, which still reflects the reasoning process and can be helpful for retrieving examples. This ensures that GraphIC is robust to inaccuracies in the generated reasoning process. Even when the reasoning process is incorrect, GraphIC still maintains high accuracy. A detailed discussion and related experiments are provided in the supplementary material.
@@ 3.2 Similarity Measure for Thought Graphs
For instance, consider a thought graph $G$ with two vertices: $v_{1}$ with attribute $x_{1}=\mathrm{Emb(\text{"add"})}$ and $v_{2}$ with $x_{2}=\mathrm{Emb(\text{"multiply"})}$, connected by an edge $v_{1}\to v_{2}$. In this setup, $W$ learns to map $x_{1}$ (”add”) to $x_{2}$ (”multiply”), indicating that ”multiply” typically follows ”add” in $G$. When this learned mapping is applied to another graph $G^{\prime}$, a high value of $\mathcal{A}(W;G^{\prime})$ suggests that $G^{\prime}$ also contains frequent ”add” $\to$ ”multiply” sequences, reflecting similar problem-solving logic. Therefore, $\mathcal{A}(W;G^{\prime})$ quantifies how well the reasoning pattern of $G$ applies to $G^{\prime}$. A higher $\mathcal{A}(W;G^{\prime})$ indicates an alignment between the two graphs’ transitions, indicating that $G^{\prime}$ follows a similar problem-solving trajectory. This metric, therefore, provides an effective measure of the similarity in reasoning between $G$ and $G^{\prime}$, making it suitable for retrieving ICEs that share compatible problem-solving logic.
@@
where $Z^{(h)}$ represents node information states at iteration $h$; $Z^{(0)}$ is initialized with node attributes $X$; II is identity matrix; $\tilde{A}=D_{A}^{-\frac{1}{2}}AD_{A}^{-\frac{1}{2}}$, $D_{A}=\mathrm{diag}(\mathrm{deg}_{\mathrm{out}}(v_{1}),\dots,\mathrm{deg}_{\mathrm{out}}(v_{1}))$; $\tilde{B}=D_{B}^{-\frac{1}{2}}BD_{B}^{-\frac{1}{2}}$, and $B$ is traceback matrix enabling backtracking to root nodes, defined as $B_{ij}=1$ if $\text{deg}_{\text{in}}(v_{j})=0$ and $\text{deg}_{\text{in}}(v_{i})>0$, otherwise $B_{ij}=0$, with $\text{deg}_{\text{in}}(v_{j})$ representing the in-degree of node $v_{j}$. A visual example of matrix $B$ is presented in the supplementary material. $\lambda$ is a balancing hyperparameter ($0\leq\lambda\leq 1$).
@@
Both theoretical analysis and empirical evidence, presented in the supplementary material, demonstrate that the rank-1 assumption has minimal impact on solving the optimization problem.
Furthermore, with the rank-1 assumption, we can obtain the closed-form solution of Equation (1), as shown in Equation (4), which greatly reduces the computational complexity of solving the optimization problem. The proof is shown in the supplementary material
@@
$\text{where}\quad U,\Sigma,V=\mathrm{SVD}(X^{\top}Z).$
@@
Secondly, our proposed similarity is asymmetric. Specifically, for two thought graphs $G_{1}$ and $G_{2}$, $s(G_{1},G_{2})\neq s(G_{2},G_{1})$. This asymmetry reflects real-world scenarios. For instance, as illustrated in the supplementary material, mathematical problem A might be a subproblem of problem B. In such a case, referencing B can help resolve A, but referencing A does not necessarily resolve B.
@@ 4.1 Experimental Setup
Datasets and Evaluations. We conduct experiments across four multi-step reasoning tasks: mathematical reasoning (GSM8K (cobbe2021training) and AQUA (ling2017program)), code generation (MBPP (austin2021program)), and logical reasoning (ProofWriter (tafjord-etal-2021-proofwriter)). Model performance is measured by answer accuracy for GSM8K, AQUA, and ProofWriter, and by the pass@1 metric (chen2021evaluating) for MBPP.
Models and Hyper-parameters. We use both an open-source and a closed-source model. For the open-source model, we select Llama-3.1-8B-Instruct (hereafter referred to as Llama-3), one of the most advanced 7B-level models. For the closed-source model, we choose GPT-4o-mini, balancing performance and testing costs. Unless explicitly mentioned otherwise, all evaluations use an 8-shot setting, which is the most common setting in chain-of-thought scenarios. We set the temperature to 1e-5 to minimize randomness, ensuring consistency across generations, and use 3 iterations in Equation (2) with $\lambda$ values from $\{0,0.1,0.2,0.3\}$ (see the supplementary material for specific details). More experiment details are in the supplementary material.
@@ 4.2 Baselines
We compare GraphIC against six training-free retrieval methods spanning random, similarity-based, diversity-based, and complexity-based approaches, including: (1) Randomrandomly selects $k$ unique ICEs from the candidate set; (2) BM25(robertson2009probabilistic)selects the top $k$ examples based on BM25 scoring; (3) BERT(devlin-etal-2019-bert)is a dense retriever using cosine similarity with BERT-base-uncased embeddings; (4) Complex-CoT(fu2022complexity)selects $k$ examples based on complexity, quantified by newline characters; (5) Auto-CoT(zhangautomatic)clusters candidates and selects the closests to each cluster center; and (6) Skill-kNN(an2023skill)prompts LLM to generate task-relevant skills for query and candidates, followed by dense retrieval.
We also compare with four training-based methods, which encompass both single-example and combination-example retrieval strategies, including: (1) EPR(rubin2022learning)is trained to retrieve the single most relevant ICE, with top $k$ examples being selected during inference; (2) CEIL(ye2023compositional)uses determinantal point processes to select ICEs balancing similarity and diversity; (3) DQ-LoRe(xiong2024dqlore)uses dual queries and low-rank approximation re-ranking to identify ICEs; and (4) GistScore(guptagistscore)encodes task-specific information into gist tokens for selecting ICEs.
@@ 4.3 Main Results
Llama-3.1
-8B-Instruct
Random 78.86 (0.87) 53.15 (1.85) 57.72 (1.06) 76.10 (2.45) 66.46 (0.84)